Water is used to extract sugar from 100,000 lb/hr
Answer #1

a)

Inlet lb/hr Stream Wet Sliced Beets Water Added Total Inlet Water Pul Sugar Total 126000 0 0 126000 174000 100000 226000 Exit

b)

% not dissolved = (12000 - 6000) /12000 * 100 = 50%

c)

% recovered = 6000/12000 * 100 = 50%